OpenStudy (anonymous):

@iGreen i will give u a meadel help!!!! rebeca had 9/10 of yard ribbon. then she gave 7/20 to her freind. how much dose she hace left. give me an explination and explain....

OpenStudy (anonymous):

9/10-7/20= Multiply the numerator and denominator by 2 of 9/10 to get to the denominator to 20 18/20-7/20=

OpenStudy (igreen):

Subtract them..first convert them into denominators of 7/20. \(\dfrac{9}{10} \rightarrow \dfrac{9 \times 2}{10 \times 2} \rightarrow \dfrac{18}{20}\) Okay, now we can subtract: \(\dfrac{18}{20} - \dfrac{7}{20}\) \(\dfrac{18-7}{20}\) Can you subtract that? @oregon8

OpenStudy (anonymous):

9

OpenStudy (igreen):

No, check again.

OpenStudy (igreen):

What's 18 - 7?

OpenStudy (anonymous):

11

OpenStudy (igreen):

Yep, that gives us \(\dfrac{11}{20}\). So Rebecca has \(\dfrac{11}{20}\) yard of ribbon left.

OpenStudy (anonymous):

thanks